OpenStudy (anonymous):

find the vertex, focus and directrix of y^2-6u+11=2x. can someone check my answers please?

OpenStudy (anonymous):

its should say Y^2-6y+11=2x

OpenStudy (mathteacher1729):

Would you know how to do it this problem if it was \[\large x^2-6x+11=2y\] instead?

OpenStudy (anonymous):

I actually can.. but i put the original problem in (y-k)^2=4p(x-h) form.. my vertex was (-1,3) and my foci were (-3/4,3) and (-5/4,3)... my directrix is x=-3/4. I'm not even sure if this is horizontal!

OpenStudy (anonymous):

@mathteacher1729 how would i know if it's horizontal or not??

OpenStudy (mathteacher1729):

The parabola is a \(\cup \) or \(\cap \) shape if it's in the form \(y = ax^2+bx+c\). It's shaped like \(\subset\) or \(\supset\) if the equation is of the form \(x = ay^2+by+c\). You can also graph it using http://www.desmos.com for free. :)

OpenStudy (mathteacher1729):

Time for me to sign off. Paul's Online Math Notes breaks this stuff down really well. Purplemath also has some good tutorials. Good luck. :)

OpenStudy (anonymous):

Thank you :)